# Leptogenesis and Neutrino Masses Via Pseudo-Dirac Gauginos

### v1

@@ I Introduction
We observe a universe that consists of only matter and virtually no antimatter. This asymmetry is often quantified through the parameter $\eta$ given by Workman et al. (2022); Aghanim et al. (2020)
@@
where $n_{B}$, $n_{\bar{B}}$ are the baryon and antibaryon number densities respectively and $n_{\gamma}$ is the photon number density. Assuming symmetric initial conditions, in order to explain why there is now a baryon asymmetry, a particle physics model is required to satisfy the following “Sakharov conditions” Sakharov (1967). 1. Baryon number violation, 2. $C$ and $CP$ violation, 3. Out of equilibrium conditions.
Baryon number is anomalously violated in the Standard Model (SM) because the $SU(2)_{L}$ interactions are chiral. At temperatures $130~{}{\rm GeV}<T<10^{12}~{}{\rm GeV}$, $B+L$ number is violated through sphaleron processes, but at low temperatures these processes are exponentially suppressed D’Onofrio et al. (2014). ($B-L$ is always conserved.) $CP$ is also violated within the SM through an irreducible phase in the Cabibo-Kobayashi-Maskawa (CKM) matrix. Although this is one of the Sakharov conditions, the SM $CP$ violation generally is not sufficient to explain the baryon asymmetry of the universe (BAU) Gavela et al. (1994); Huet and Sather (1995) [^1] Recently a BSM model that uses the observed $CP$ violation in the SM was introduced Elor et al. (2024).. Finally, the SM does not satisfy the out-of-equilibrium conditions. Hence beyond-the-SM (BSM) physics is needed to explain the BAU.
In addition to failing to explain the BAU, the SM cannot explain the origin of neutrino masses either. The fact that neutrinos oscillate between their flavor eigenstates requires that they have masses. Neutrino oscillation data give us two mass differences Esteban et al. (2024):
@@
Note that the mass ordering and the absolute masses are currently not known and the lightest neutrino is allowed to be massless. In order to explain how neutrinos acquire their masses, new physics (NP) is needed. One of the most studied NP mechanisms to explain the origin of neutrino masses is the seesaw mechanism with a heavy right-handed (RH) Majorana singlet fermion (see, e.g., King (2016); Gavela et al. (2009); Zhi-zhong Xing and Zhen-hua Zhao (2021) and the references therein). In this scenario, the light neutrino masses are proportional to $\sim v^{2}/M$, where $v$ is the Higgs vacuum expectation value (vev) and $M$ is the Majorana mass of the RH neutrino. To match the mass scale of the SM neutrinos, the seesaw scale usually needs to be very high, $M\sim 10^{14}~{}{\rm GeV}$. Alternatively, if RH neutrinos are pseudo-Dirac fermions, having both Dirac and Majorana masses, neutrino masses are generated via the inverse-seesaw (ISS) mechanism Wyler and Wolfenstein (1983); Mohapatra (1986); Mohapatra and Valle (1986). In ISS, light neutrino masses are instead proportional to a small Majorana mass that slightly breaks the lepton number.
@@
A common approach to solving the origins of the BAU and the neutrino masses is leptogenesis Fukugita and Yanagida (1986). (See Davidson et al. (2008) for an extensive review.) In leptogenesis, a heavy RH Majorana neutrino decays out-of-equilibrium, producing a lepton asymmetry. The asymmetry in leptons is then partially converted into the baryon asymmetry through sphalerons. Minimal leptogenesis models require at least two RH neutrinos with masses $\mathcal{O}(10^{10-12}~{}{\rm GeV})$, which are impossible to probe in particle physics experiments. In models such as resonant leptogenesis Pilaftsis and Underwood (2004); Pilaftsis (2005); Pilaftsis and Underwood (2005); Chauhan and Dev (2023); Klarić et al. (2021); King et al. (2024); Das and Orikasa (2024) and ARS leptogenesis Akhmedov et al. (1998); Drewes et al. (2018, 2022); Caputo et al. (2019); Baumholzer et al. (2018), the Majorana neutrino mass scale could be much lower, possibly $\mathcal{O}({\rm TeV})$. Generally, this lowering of the mass scale requires almost-degenerate RH neutrinos, resulting in an enhancement of the lepton asymmetry by the small mass difference between the RH neutrinos.
@@
In this work, we analyze a leptogenesis scenario within a neutrino mass mechanism with pseudo-Dirac bino and wino Coloma and Ipek (2016); Ayber and Ipek (2023). Psuedo-Dirac gauginos are naturally expected in $U(1)_{R}-$symmetric minimally supersymmetric SM (MSSM), in which supersymmetric partners are charged under a global $U(1)_{R}$ symmetry. In Coloma and Ipek (2016), it was shown that a pseudo-Dirac bino can act like a (pseudo-Dirac) RH neutrino, generating light neutrino masses through an ISS mechanism. In Ayber and Ipek (2023), it was further shown that the pseudo-Dirac wino in the same model could also be involved in neutrino mass generation, giving rise to a hybrid type I+III ISS scenario with rich phenomenology.
@@
A pseudo-Dirac fermion can have particle–antiparticle oscillations, much like the neutral meson oscillations in the SM. Furthermore, if the oscillation frequency and the decay width of the particles are comparable, the effects of possible $CP$ violation in the system are enhanced Ipek et al. (2014). In Ipek and March-Russell (2016), such $CP$ violation via oscillations was shown to result in successful baryogenesis, given that the pseudo-Dirac fermion has $B$-violating decays. In the same work, a pseudo-Dirac bino in a generic $U(1)_{R}-$symmetric MSSM was discussed as a specific example, focusing on baryon number violation via $R-$parity violating interactions for a benchmark bino mass of $300$ GeV.
@@
In this work, we show that the hybrid bino-wino scenario in Ayber and Ipek (2023) can lead to successful leptogenesis via bino–antibino oscillations in the early universe, explaining both the neutrino masses and the BAU. We explore the parameter space for successful leptogenesis, analyzing the bino mass between 500 GeV and 5 TeV, and the sfermion masses $\mathcal{O}(10-100~{}{\rm TeV})$.
@@ II The Model
In this section, we summarize the neutrino mass mechanism in a $U(1)_{R-L}$–symmetric SUSY model introduced in Coloma and Ipek (2016) and extended in Ayber and Ipek (2023). For a viable leptogenesis scenario, we focus on Ayber and Ipek (2023), which we will motivate shortly.
In MSSM with a global $U(1)_{R}$ symmetry, superpartners have $+1$ $R-$charges, while SM fields are not charged under $U(1)_{R}$. Due to this global charge, gauginos cannot have Majorana masses. In order to give mass to gauginos, adjoint Dirac partners with $-1$ $R-$charges are introduced. (See Table 1 for partial field content and charge assignments.) The relevant Dirac partners for this work will be the singlino $S$ and tripletino $T$ which are Dirac partners for bino and wino, respectively.
@@
Table 1: A partial superfield content of the model and their charge assignments. The $R-L$ charge is defined as the $U(1)_{R}$-charge minus the lepton number of the field. $L_{i},\ E_{i}^{c}$ are the lepton superfields where the subindex i indicates the fermion generation. The fermionic components of the superfields $R_{u,d}$ are the Dirac partners of the Higgsinos $H_{u,d}$. $\Phi_{S},\Phi_{T}$ are the superfields that have the same SM charges as $W^{\alpha}_{\tilde{B}}$ and $W^{\alpha}_{\widetilde{W}}$. Their fermionic components, $S$ and $T$ are the Dirac partners of the bino and the wino, respectively.
Dirac gaugino masses can be generated via the soft terms Fox et al. (2002)
@@
Although we started with a $U(1)$ symmetry, it must be broken due to gravity like any other global symmetry. At the end, Majorana masses for the gauginos will be generated via anomaly mediation Randall and Sundrum (1999); Giudice et al. (1998); Gherghetta et al. (1999),
@@
where $\beta(g_{\tilde{\chi}})$ is the beta function of the relevant gaugino and $F_{\phi}$ is a conformal factor satisfying
@@
In Ayber and Ipek (2023), the global symmetry was extended to a $U(1)_{R-L}$ symmetry, where $L$ is the lepton number. The resulting pseudo-Dirac electroweakino fields can act as right-handed neutrinos and generate non-zero neutrino masses when the following effective operators are considered:
@@
where $\phi=1+\theta^{2}F_{\phi}$ is the conformal compensator and $f^{i}_{\tilde{\chi}},d^{i}_{\eta}$, for $\tilde{\chi}=\tilde{B},\tilde{W},\eta=S,T$, and $i=e,\mu,\tau$ are dimensionless, complex coefficients. Note that in Ayber and Ipek (2023), a scenario with $F_{\phi}=m_{3/2}$ was considered, which required $(m_{S}+m_{T})\sim m_{3/2}\sim O(100~{}{\rm keV})$ for $\Lambda_{M}\sim O(500~{}{\rm TeV})$ to generate the correct neutrino mass scale. Later we will see that for leptogenesis, we need a small mass splitting for the bino mass eigenstates, $m_{\tilde{B}}+m_{S}\sim F_{\phi}\sim(10^{-3}-10)$ eV, and a messenger scale $\Lambda_{M}\gtrsim 10^{7}$ TeV. Such a small $F_{\phi}$ with a large messenger scale would make neutrino masses far too small to match the neutrino oscillation data. Hence, we allow instead for the following operator to appear,
@@
where $Y^{i}_{\tilde{B},\tilde{W}}=f^{i}_{\tilde{B},\tilde{W}}M_{\tilde{B},\tilde{W}}/\Lambda_{M}$ and $G^{i}_{S,T}=d^{i}_{S,T}\tilde{m}/\Lambda_{M}$. The mass matrix above has a hybrid type-I+III ISS structure, and in its most general form, it can generate all non-zero neutrino mass eigenvalues. In Ayber and Ipek (2023), a simplified scenario with $\mathbf{G}_{S,T}=0$ was analyzed. The resulting light neutrino masses for normal ordering were determined to be
@@
where $m_{2}<m_{3}$. The mass-squared splitting ratios $\Delta m^{2}_{\rm solar}/\Delta m^{2}_{\rm atm}$, sets the parameter $\beta\simeq 0.13$. Notably, the light neutrino masses are independent of the Dirac masses.
@@
The out-of-equilibrium condition can be met for $\Gamma_{\tilde{B}}<H(T=M_{\tilde{B}})$ where $H(T)$ is the Hubble expansion rate given by,
@@
Figure 1: Regions of bino decay widths relevant for leptogenesis with respect to the bino mass $M_{\tilde{B}}$ and the messenger scale $\Lambda_{M}$. Here we take $\tilde{m}=0.2M_{\tilde{B}}$. In the bottom purple region bino decays at a temperature above its mass, $T>M_{\tilde{B}}$. In the upper green region, bino lives long enough to decay below $T_{\rm EW}\simeq 130$ GeV, when sphalerons are no longer active.
Thus, we consider a bino heavier than $500\rm~{}GeV$ and messenger scales $\Lambda_{M}\gtrsim\mathcal{O}(10^{7}~{}\rm TeV)$. (See Figure 1.) Furthermore, the amount of $CP$ violation needed to generate the correct baryon asymmetry will be selected for a (wide) range of $m_{\tilde{B}}+m_{S}$ and $\tilde{m}$. (See fig. 3.) We also note that the condition on the lifetime is the reason why the pure-bino case in Coloma and Ipek (2016) cannot account for both neutrino masses and the BAU. The parameters that allow out-of-equilibrium bino decays result in a too small contribution to neutrino masses. In the bino-wino hybrid scenario of Ayber and Ipek (2023) there is enough freedom to accommodate both the light neutrino masses and a long-lived bino.
@@ III Bino–Antibino Oscillations in the Early Universe
Pseudo-Dirac gaugino oscillations in vacuum at zero temperature were studied in the literature Grossman et al. (2013); Ipek et al. (2014). In Ipek et al. (2014) it was shown that $CP$ violation in pseudo-Dirac gluino decays can lead to interesting same-sign dilepton signatures at the LHC. In the early universe, these oscillations would take place in a hot and dense medium in which the pseudo-Dirac binos interact efficiently. Furthermore, the expansion of the universe must be taken into account. Using the density matrix formalism Tulin et al. (2012), such an analysis was carried out in Ipek and March-Russell (2016) for a baryogenesis scenario with $R$-parity violating bino interactions. (See also Chowdhury (2024) for an expansion of that work.) Here we focus on a leptogenesis scenario in which the $L$-violating interactions also explain the origin of neutrino masses. In this section, we will briefly lay out the Boltzmann equations that govern the CP-violating pseudo-Dirac bino oscillations and decays in the early universe.
@@ III.1 Bino oscillations
Before moving into the early universe, let us first summarize the oscillations of pseudo-Dirac binos, a phenomenon very similar to neutral-meson oscillations. We will be working with the Dirac field and its charge conjugate
@@
$\mathbf{H}=\mathbb{M}-\frac{i}{2}\mathbf{\Gamma}\,,$ (13)
$=\begin{pmatrix}M_{D}&m\\ m^{*}&M_{D}\end{pmatrix}-\frac{i}{2}\frac{M_{\tilde{B}}}{8\pi}\begin{pmatrix}|g_{\tilde{B}}|^{2}+|g_{S}|^{2}&2g_{\tilde{B}}^{*}g_{S}\\ 2g_{\tilde{B}}g_{S}^{*}&|g_{\tilde{B}}|^{2}+|g_{S}|^{2}\end{pmatrix}\,,$
where $M_{D}\simeq M_{\tilde{B}}$ and $m\simeq(m_{\tilde{B}}+m_{S}^{\ast})/2$ are the renormalized Dirac and Majorana masses [^2] The Dirac mass is multiplicatively renormalized while the Majorana mass will get contributions from interactions. The loop contributions to the Majorana mass are $\delta\sim\frac{|g_{\tilde{B}}g_{S}^{\ast}|}{(4\pi)^{2}}M_{D}\sim\frac{M_{\tilde{B}}^{3}}{(4\pi)^{2}\Lambda_{M}^{2}}\lesssim O(10^{-14}~{}{\rm eV})$., respectively, and $g_{\tilde{B}}=\sum_{i}f_{i}M_{\tilde{B}}/\Lambda_{M}$ and $g_{S}=\sum_{i}d_{i}\tilde{m}/\Lambda_{M}$. (We have taken the Dirac mass $M_{\tilde{B}}$ to be real.) As discussed before, we will consider binos that are heavier than $\sim 500$ GeV, and thus they can decay into final states with a Higgs boson, gauge bosons, and leptons via their mixing with the neutrinos. For the lepton asymmetry, we will only track the decays of $\tilde{B},S\to h\ell$.
@@
Although both the mass and the width part of the Hamiltonian in eq. 13 can be complex, there is only one physical phase in this system. We choose a real mass matrix, putting the $CP$ violating phase in $\mathbf{\Gamma}$. The decay matrix can be re-written as
$\mathbb{\Gamma}=\Gamma_{\tilde{B}}\begin{pmatrix}1&2re^{i\phi}\\ 2re^{-i\phi}&1\end{pmatrix}\,,$ (17)
@@
In the leptogenesis section, we will set $\tilde{m}\simeq 0.2M_{\tilde{B}}$, i.e. $r\simeq 0.2$ as a benchmark value. This value results in a large $CP$ violation but we will show that the observed BAU can be generated with smaller values, $r\sim 10^{-4}$, as well.
The eigenstates of the Hamiltonian are a superposition of bino and antibino states. A purely particle state at $t=0$ will evolve to acquire an antiparticle component and vice versa for an initial antiparticle state. The time evolution of an initial particle or antiparticle state is given by
$\begin{split}|\psi_{\tilde{B}}(t)\rangle&=g_{+}(t)|\psi_{\tilde{B}}\rangle-\frac{q}{p}g_{-}(t)|\psi_{\tilde{B}}^{C}\rangle\,,\\ |\psi_{\tilde{B}}^{c}(t)\rangle&=g_{+}(t)|\psi_{\tilde{B}}^{C}\rangle-\frac{p}{q}g_{-}(t)|\psi_{\tilde{B}}\rangle\,,\end{split}$ (19)

### v2

@@ I Introduction
We observe a universe that consists of only matter and virtually no antimatter. This asymmetry is often quantified through the parameter $\eta$ given by ParticleDataGroup:2022pth; Planck:2018vyg
@@
where $n_{B}$, $n_{\bar{B}}$ are the baryon and antibaryon number densities respectively and $n_{\gamma}$ is the photon number density. Assuming symmetric initial conditions, in order to explain why there is now a baryon asymmetry, a particle physics model is required to satisfy the following “Sakharov conditions” Sakharov:1967dj. 1. Baryon number violation, 2. $C$ and $CP$ violation, 3. Out of equilibrium conditions.
Baryon number is anomalously violated in the Standard Model (SM) because the $SU(2)_{L}$ interactions are chiral. At temperatures $130~{\rm GeV}<T<10^{12}~{\rm GeV}$, $B+L$ number is violated through sphaleron processes, but at low temperatures these processes are exponentially suppressed PhysRevLett.113.141602. (BLB-L is always conserved.) $CP$ is also violated within the SM through an irreducible phase in the Cabibo-Kobayashi-Maskawa (CKM) matrix. Although this is one of the Sakharov conditions, the SM $CP$ violation generally is not sufficient to explain the baryon asymmetry of the universe (BAU) Gavela:1993ts; Huet:1994jb [^1] Recently, a BSM model that uses the observed $CP$ violation in the SM was introduced Elor:2024cea.. Finally, the SM does not satisfy the out-of-equilibrium conditions. Hence, beyond-the-Standard Model (BSM) physics is needed to explain the BAU.
In addition to failing to explain the BAU, the SM cannot explain the origin of neutrino masses either. The fact that neutrinos oscillate between their flavor eigenstates requires that they have masses. Neutrino oscillation data give us two mass differences Esteban:2024eli:
@@
Note that currently, the mass ordering and the absolute masses are not known, and the lightest neutrino is allowed to be massless. In order to explain how neutrinos acquire their masses, new physics (NP) is required. One of the most studied NP mechanisms to explain the origin of neutrino masses is the seesaw mechanism with a heavy right-handed (RH) Majorana singlet fermion (see, e.g., KING:2016456; Gavela:2009cd; Xing:2021 and the references therein). In this scenario, the light neutrino masses are proportional to $\sim v^{2}/M$, where $v$ is the Higgs vacuum expectation value (vev) and $M$ is the Majorana mass of the RH neutrino. To match the mass scale of the SM neutrinos, the seesaw scale usually needs to be very high, $M\sim 10^{14}~{\rm GeV}$. Alternatively, if RH neutrinos are pseudo-Dirac fermions, having both Dirac and Majorana masses, neutrino masses are generated via the inverse-seesaw (ISS) mechanism WYLER:1983205; Mohapatra:1986aw; Mohapatra:1986bd. In ISS, light neutrino masses are instead proportional to a small Majorana mass that slightly breaks the lepton number.
@@
A common approach to solving the origins of the BAU and the neutrino masses is leptogenesis Fukugita:1986hr. (See Davidson_2008 for an extensive review.) In leptogenesis, a heavy RH Majorana neutrino decays out-of-equilibrium, producing a lepton asymmetry. The asymmetry in leptons is then partially converted into the baryon asymmetry through sphalerons. Minimal leptogenesis models require at least two RH neutrinos with masses $\mathcal{O}(10^{10-12}~{\rm GeV})$, which are impossible to probe in particle physics experiments. In models such as resonant leptogenesis Pilaftsis:2003gt; Pilaftsis:2004xx; Pilaftsis:2005rv; Chauhan:2021xus; Klaric:2021cpi; King:2024idj; Das:2024gua and ARS leptogenesis Akhmedov:1998qx; Drewes:2017zyw; Drewes:2021nqr; Caputo:2018zky; Baumholzer:2018sfb, the Majorana neutrino mass scale could be much lower, possibly $\mathcal{O}({\rm TeV})$. Generally, this lowering of the mass scale requires almost-degenerate RH neutrinos, resulting in an enhancement of the lepton asymmetry by the small mass difference between the RH neutrinos.
@@
In this work, we analyze a leptogenesis scenario within a neutrino mass mechanism with pseudo-Dirac bino and wino Coloma:2016vod; Ayber:2023diw. Pseudo-Dirac gauginos are naturally expected in $U(1)_{R}$-symmetric minimally supersymmetric SM (MSSM), in which supersymmetric partners are charged under a global $U(1)_{R}$ symmetry. In Coloma:2016vod, it was shown that a pseudo-Dirac bino can act like a (pseudo-Dirac) RH neutrino, generating light neutrino masses through an ISS mechanism. In Ayber:2023diw, it was further shown that the pseudo-Dirac wino in the same model could also be involved in neutrino mass generation, giving rise to a hybrid type I+III ISS scenario with rich phenomenology.
@@
A pseudo-Dirac fermion can have particle–antiparticle oscillations, much like the neutral meson oscillations in the SM. Furthermore, if the oscillation frequency and the decay width of the particles are comparable, the effects of possible $CP$ violation in the system are enhanced Ipek:2014moa. In Ipek_2016, such $CP$ violation via oscillations was shown to result in successful baryogenesis, given that the pseudo-Dirac fermion has $B$-violating decays. In the same work, a pseudo-Dirac bino in a generic $U(1)_{R}$-symmetric MSSM was discussed as a specific example, focusing on baryon number violation via $R$-parity violating interactions for a benchmark bino mass of $300$ GeV.
@@
In this work, we show that the hybrid bino-wino scenario in Ayber:2023diw can lead to successful leptogenesis via bino–antibino oscillations in the early universe, explaining both the neutrino masses and the BAU. We explore the parameter space for successful leptogenesis, analyzing the bino mass between 500 GeV and 5 TeV, and the sfermion masses $\mathcal{O}(10-100~{\rm TeV})$.
@@ II The Model
In this section, we summarize the neutrino mass mechanism in a $U(1)_{R-L}$–symmetric SUSY model introduced in Coloma:2016vod and extended in Ayber:2023diw. For a viable leptogenesis scenario, we focus on Ayber:2023diw, which we will motivate shortly.
In MSSM with a global $U(1)_{R}$ symmetry, superpartners have +1+1 $R$-charges, while SM fields are not charged under $U(1)_{R}$. Due to this global charge, gauginos cannot have Majorana masses. In order to give mass to gauginos, adjoint Dirac partners with 1-1 $R$-charges are introduced. (See Table 1 for partial field content and charge assignments.) The relevant Dirac partners for this work will be the singlino $S$ and tripletino $T$, which are Dirac partners for bino and wino, respectively.
@@
Table 1: A partial superfield content of the model and their charge assignments. The RLR-L charge is defined as the $U(1)_{R}$-charge minus the lepton number of the field. $L_{i},\ E_{i}^{c}$ are the lepton superfields where the subindex i indicates the fermion generation. The fermionic components of the superfields $R_{u,d}$ are the Dirac partners of the Higgsinos $H_{u,d}$. $\Phi_{S},\Phi_{T}$ are the superfields that have the same SM charges as $W^{\alpha}_{\tilde{B}}$ and $W^{\alpha}_{\widetilde{W}}$. Their fermionic components, $S$ and $T$, are the Dirac partners of the bino and the wino, respectively.
Dirac gaugino masses can be generated via the soft terms Fox:2002bu
@@
Although we started with a $U(1)$ symmetry, it must be broken due to gravity, like any other global symmetry. At the end, Majorana masses for gauginos will be generated via anomaly mediation Randall:1998uk; Giudice:1998xp; Gherghetta:1999am,
@@
[…]
@@
In Ayber:2023diw, the global symmetry was extended to a $U(1)_{R-L}$ symmetry, where $L$ is the lepton number. The resulting pseudo-Dirac electroweakino fields can act as right-handed neutrinos and generate non-zero neutrino masses when the following effective operators are considered:
@@
where $\phi=1+\theta^{2}F_{\phi}$ is the conformal compensator and $f^{i}_{\tilde{\chi}},d^{i}_{\eta}$, for $\tilde{\chi}=\tilde{B},\tilde{W},\eta=S,T$, and $i=e,\mu,\tau$ are dimensionless, complex coefficients. Note that in Ayber:2023diw, a scenario with $F_{\phi}=m_{3/2}$ was considered, which required $(m_{S}+m_{T})\sim m_{3/2}\sim\mathcal{O}(100~{\rm keV})$ for $\Lambda_{M}\sim\mathcal{O}(500~{\rm TeV})$ to generate the correct neutrino mass scale. Later, we will see that for leptogenesis, we need a small mass splitting for the bino mass eigenstates, $m_{\tilde{B}}+m_{S}\sim F_{\phi}\sim(10^{-3}-10)$ eV, and a messenger scale $\Lambda_{M}\gtrsim 10^{7}$ TeV. Such a small $F_{\phi}$ with a large messenger scale would make neutrino masses far too small to match the neutrino oscillation data. Hence, we allow instead for the following operator to appear,
@@
where $Y^{i}_{\tilde{B},\tilde{W}}=f^{i}_{\tilde{B},\tilde{W}}M_{\tilde{B},\tilde{W}}/\Lambda_{M}$ and $G^{i}_{S,T}=d^{i}_{S,T}\tilde{m}/\Lambda_{M}$. The mass matrix above has a hybrid type-I+III ISS structure, and in its most general form, it can generate all non-zero neutrino mass eigenvalues. In Ayber:2023diw, a simplified scenario with $\mathbf{G}_{S,T}=0$ was analyzed. The resulting light neutrino masses for normal ordering were determined to be
@@
where $m_{2}<m_{3}$. The mass-squared splitting ratios $\Delta m^{2}_{\rm solar}/\Delta m^{2}_{\rm atm}$, set the parameter $\beta\simeq 0.13$. Notably, the light neutrino masses are independent of the Dirac masses.
@@
[…]
@@
Figure 1: Regions of bino decay widths relevant for leptogenesis with respect to the bino mass $M_{\tilde{B}}$ and the messenger scale $\Lambda_{M}$. Here we take $\tilde{m}=0.2M_{\tilde{B}}$. In the bottom purple region, bino decays at a temperature above its mass, $T>M_{\tilde{B}}$. In the upper green region, bino lives long enough to decay below $T_{\rm EW}\simeq 130$ GeV, when sphalerons are no longer active.
Thus, we consider a bino heavier than $500\rm~GeV$ and messenger scales $\Lambda_{M}\gtrsim\mathcal{O}(10^{7}~\rm TeV)$. (See Figure 1.) Furthermore, the amount of $CP$ violation needed to generate the correct baryon asymmetry will be selected for a (wide) range of $m_{\tilde{B}}+m_{S}$ and $\tilde{m}$. (See fig. 3.) We also note that the condition on the lifetime is the reason why the pure-bino case in Coloma:2016vod cannot account for both neutrino masses and the BAU. The parameters that allow for out-of-equilibrium bino decays result in a too small contribution to neutrino masses. In the bino-wino hybrid scenario of Ayber:2023diw, there is enough freedom to accommodate both light neutrino masses and a long-lived bino.
@@ III Bino–Antibino Oscillations in the Early Universe
Pseudo-Dirac gaugino oscillations in vacuum at zero temperature were studied in the literature Grossman:2012nn; Ipek:2014moa. In Ipek:2014moa, it was shown that $CP$ violation in pseudo-Dirac gluino decays can lead to interesting same-sign dilepton signatures at the LHC. In the early universe, these oscillations would take place in a hot, dense medium in which the pseudo-Dirac binos interact efficiently. Furthermore, the expansion of the universe must be taken into account. Using the density matrix formalism Tulin:2012re, such an analysis was carried out in Ipek_2016 for a baryogenesis scenario with $R$-parity violating bino interactions. (See also Chowdhury:2024lxx for an expansion of that work.) Here we focus on a leptogenesis scenario in which the $L$-violating interactions also explain the origin of neutrino masses. In this section, we will briefly lay out the Boltzmann equations that govern the CP-violating pseudo-Dirac bino oscillations and decays in the early universe.
@@ III.1 Bino oscillations
Before moving on to the early universe, let us summarize the oscillations of pseudo-Dirac binos, a phenomenon closely analogous to neutral-meson oscillations. We will be working with the Dirac field and its charge conjugate
@@
$\mathbf{H}$ $=\textbf{M}-\frac{i}{2}\mathbf{\Gamma}\,$, (13)
$=\begin{pmatrix}M_{D}&\hskip-4.26773ptm\\ m^{*}&\hskip-4.26773ptM_{D}\end{pmatrix}-\frac{i}{2}\frac{M_{\tilde{B}}}{8\pi}\begin{pmatrix}|g_{\tilde{B}}|^{2}+|g_{S}|^{2}&\hskip-4.26773pt2g_{\tilde{B}}^{*}g_{S}\\ 2g_{\tilde{B}}g_{S}^{*}&\hskip-4.26773pt|g_{\tilde{B}}|^{2}+|g_{S}|^{2}\end{pmatrix}$,
where $M_{D}\simeq M_{\tilde{B}}$ and $m\simeq(m_{\tilde{B}}+m_{S}^{\ast})/2$ are the renormalized Dirac and Majorana masses [^2] The Dirac mass is multiplicatively renormalized while the Majorana mass will get contributions from interactions. The loop contributions to the Majorana mass are $\delta\sim\frac{|g_{\tilde{B}}g_{S}^{\ast}|}{(4\pi)^{2}}M_{D}\sim\frac{M_{\tilde{B}}^{3}}{(4\pi)^{2}\Lambda_{M}^{2}}\lesssim\mathcal{O}(10^{-14}~{\rm eV})$., respectively, and $g_{\tilde{B}}=\sum_{i}f_{i}M_{\tilde{B}}/\Lambda_{M}$ and $g_{S}=\sum_{i}d_{i}\tilde{m}/\Lambda_{M}$. (We have taken the Dirac mass $M_{\tilde{B}}$ to be real.) As discussed before, we will consider binos that are heavier than $\sim 500$ GeV, and thus they can decay into final states with a Higgs boson, gauge bosons, and leptons via their mixing with the neutrinos. For the lepton asymmetry, we will only track the decays of $\tilde{B},S\to h\ell$.
@@
Although both the mass and the width part of the Hamiltonian in eq. 13 can be complex, there is only one physical phase in this system. We choose a real mass matrix, putting the $CP$ violating phase in $\mathbf{\Gamma}$. The decay matrix can be rewritten as
$\mathbf{\Gamma}=\Gamma_{\tilde{B}}\begin{pmatrix}1&2re^{i\phi}\\ 2re^{-i\phi}&1\end{pmatrix}\,$, (17)
@@
In the leptogenesis section, we will set $\tilde{m}\simeq 0.2M_{\tilde{B}}$, i.e. $r\simeq 0.2$ as a benchmark value. This choice leads to a large $CP$ violation, but we will show that the observed BAU can also be generated with smaller values, $r\sim 10^{-4}$.
The eigenstates of the Hamiltonian are a superposition of bino and antibino states. A purely particle state at $t=0$ will evolve to acquire an antiparticle component, and vice versa for an initial antiparticle state. The time evolution of an initial particle or antiparticle state is given by
[…]
@@
$N_{\ell}=\int_{0}^{\infty}dt\frac{d\Gamma(\psi_{\tilde{B}}\to h\ell)}{dt}\,,~\bar{N}_{\ell}=\int_{0}^{\infty}dt\frac{d\Gamma(\psi_{\tilde{B}}^{C}\to h\ell)}{dt}\,$,
and $\Gamma(\psi_{\tilde{B}}\to h\ell)$ is the partial decay width for an initial $|\psi_{\tilde{B}}\rangle$ state to decay into a lepton state. Note that the above measure of asymmetry is time-integrated. Setting $\Delta\Gamma/(2\Gamma_{\tilde{B}})=0$, in the limit $r\ll 1$ and $m\gg r\Gamma_{\tilde{B}}$, we have the following approximation,
where $x\equiv\Delta m/\Gamma_{\tilde{B}}\simeq 2m/\Gamma_{\tilde{B}}$. As expected, the $CP$ violation is maximized for $x\sim 1$. However, the parameter regime we consider will be where $x\simeq\mathcal{O}(1-1000)$, which still results in enough baryon asymmetry. Since this approximation is valid in the parameter regime we consider, we will use eq. 22 in the calculation of the baryon asymmetry later in eq. 35.
III.2 Other bino interactions
Unlike generic RH neutrinos in various seesaw models, the bino in this model has copious interactions with the SM particles. Elastic scatterings of the bino with the SM fermions in the plasma and bino–antibino annihilations affect the bino–antibino oscillations in the early universe, and hence the baryon asymmetry that will be generated through their decays. Assuming a common sfermion mass, $M_{\rm sf}$, the relevant effective interaction Lagrangian, after integrating out heavy sfermions, can be written as Hsieh:2007wq,